### v2

# Relativistic stars in $f(Q)$-gravity: Exact analytic solution for the power-law case $f(Q)=Q+b\>Q^{\nu}$

Abstract
We investigate static spherically symmetric spacetimes within the framework of symmetric teleparallel $f(Q)$ gravity in order to describe relativistic stars. We adopt a specific ansatz for the background geometry corresponding to a singularity-free space-time. We obtain an expression for the connection, which allows the derivation of solutions for any $f(Q)$ theory in this context. Our approach aims to address a recurring error appearing in the literature, where even when a connection compatible with spherical symmetry is adopted, the field equation for the connection is systematically omitted and not checked if it is satisfied. For the stellar configuration, we concentrate on the power-law model $f(Q)=Q+\alpha Q_{0}\left(\frac{Q}{Q_{0}}\right)^{\nu}$. The de Sitter-Schwarzschild geometry naturally emerges as an attractor beyond a certain radius, we thus utilize it as the external solution beyond the boundary of the star. We perform a detailed investigation of the physical characteristics of the interior solution, explicitly determining the mass function, analyzing the resulting gravitational fluid properties and deriving the angular and radial speed of sound.
Stars; Static Spherically Symmetric; Symmetric teleparallel $f(Q)$-gravity
I Introduction
The analysis of cosmological observations [1, 2, 3, 4, 5, 6, 7, 8] challenges General Relativity (GR) as a gravitational theory on very large scales. Modifying the Einstein-Hilbert action by introducing geometric scalars, opens new directions in the explanation of gravitational phenomena [9, 10, 11]. The resulting modified field equations aim to provide corrections to GR, although the deviations from the classical predictions need to be suppressed at scales relevant to laboratory experiments or solar system tests. The rapid progress in observational cosmology and multimessenger astrophysics, allows now for precision tests of fundamental physics, both on the scale of the observable Universe, and in the strong gravity regime of relativistic compact objects.
Possible quantum corrections to GR are sometiems expressed through modifications of the Einstein-Hilbert Action by the addition of higher-order curvature invariants [12, 13, 14, 15]. The Starobinsky model of inflation [16] belongs to this family of theories; an $R^{2}$ term is introduced in order to drive the dynamics that explain the early-time acceleration of the universe. The general family of modified theories in which the Starobinsky model belongs to is the $f\left(R\right)$-gravity [17]. These theories are also related, through conformal mappings, to scalar tensor theories like e.g. the Brans-Dicke gravity [18].
Recently, Symmetric Teleparallel Equivalent General Relativity (STEGR) [19] and its extensions [20, 21, 22, 23, 24, 25, 26, 27] have drawn the attention of the academic community. For an extended review, we refer the interested reader to [28]. In STEGR, the manifold is equipped with a metric tensor and a symmetric, flat connection. From these properties of the connection, only the nonmetricity tensor $Q_{\kappa\mu\nu}$ survives, and the ensuing nonmetricity scalar, $Q$, is the fundamental element of the Lagrangian density of the gravitational Action Integral. The $Q$ is defined so that its difference from the Ricci scalar, calculated with the Levi-Civita connection of the metric tensor, is just a boundary term [28]. As a result the Action Integral of STEGR leads to gravitational equations equivalent to those of GR. However, this equivalence is fragile, as it is not carried over when one considers Lagrangians with a nonlinear dependence on the relative scalars.
The $f\left(Q\right)$-theory of gravity, which has been introduced recently and aims to describe theories with a nonlinear dependence on $Q$, has been widely applied to describe the cosmic acceleration, including dark energy. For relevant references see [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44] and references therein. At this point, it is important to mention that, while $f\left(Q\right)$-gravity suffers from the appearance of ghosts, or strong coupling in cosmological perturbations [45, 46], it nevertheless serves as a case study to understand the effects of the connection in gravity. The choice of connection in $f\left(Q\right)$-gravity is not always unique, meaning that different connections lead to different gravitational models.
There are several studies in the literature that investigate static, spherically symmetric spacetimes in $f\left(Q\right)$-gravity. Exact solutions describing black holes have been presented in [47, 48, 49, 50, 51, 52], while black holes in nonmetricity scalar-tensor theories were studied in [50, 53]. The latter, is a more general theory which includes $f\left(Q\right)$-gravity as a limiting case. The setting is similar to that which relates Brans-Dicke to $f\left(R\right)$-gravity. In [53], an interesting result was found: namely, there are no static spherically symmetric black hole solutions in $f\left(Q\right)$-gravity when the two metric coefficients, $g_{tt}$ and $g_{rr}$, are constrained by $g_{rr}g_{tt}=1$. Moreover, the use of the coincidence gauge for static, spherically symmetric spacetimes was discussed in [54], and possible constraints from solar system tests in the theory are discussed in [55].
As has been mentioned in [56], the timeline of the investigation of the properties of relativistic and compact stars, initiated in 1916 with Schwarzschild’s solution to Einstein’s field equations; an analysis that has been fundamental to astrophysics. The solution may be interpreted as the gravitational field generated by a body characterized by spherical symmetry [57]. That work boosted subsequent studies on stellar modelling and internal composition of astronomical objects characterized by ultra dense matter content. Under those extreme conditions that cannot reproduced on terrestrial experiments, the Newtonian description becomes inadequate, and thus a full general relativistic treatment is required. Thanks to their unique properties, relativistic stars comprise ideal cosmic laboratories capable of constraining new physics and alternative theories of gravity.
When studying stars the authors usually focus their attention on astrophysical objects made of matter viewed as a perfect fluid, where there is a unique pressure along all spatial dimensions, $p_{r}=p_{\theta}=p_{\phi}$. However, celestial bodies are not necessarily made of isotropic fluids. In fact, under certain conditions matter content of compact objects may become anisotropic. The review article of Ruderman [58] long time ago mentioned such a possibility: there the author makes for the first time the observation that in very dense media anisotropies can be generated by interactions between relativistic particles. A comprehensive study of fluid spheres made of anisotropic matter was presented in [59] in Newtonian as well as relativistic gravity. In the context of Einstein’s General Relativity, anisotropic fluids have been extensively studied, with particular emphasis on stellar modelling and their impact on structural properties of compact objects [60, 61].
In spite of many successful predictions of GR and its mathematical elegance, in most of the cases of interest the study and analysis of realistic physical situations comprises a formidable task, due to the fact that Einstein’s field equations are highly non-linear coupled partial differential equations. Therefore, finding exact analytic solutions to the field equations of Einstein’s gravity has always been an interesting and challenging topic, keeping researchers busy for decades. For exact known solutions to Einstein’s field equations, see [62]. Nowadays there is a number of approaches that allow us to obtain exact analytic solutions to the gravitational field equations, such as the Minimal Geometric Deformation [63], the vanishing complexity factor formalism [64] as well as the Karmarkar condition in gravity [65].
In this work, we focus on analyzing interior static spherically symmetric solutions within the framework of $f\left(Q\right)$-gravity. To realize this investigation, we consider a specific ansatz for the geometry and examine the physical properties of the gravitational fluid. We derive a solution for the connection, which provides a general setting that can be applied to any $f(Q)$ theory in the search of stellar solutions. We make use of the relevant solution to investigate stellar configurations in a power-law theory of the form $f(Q)=Q+\alpha\,Q_{0}\left(\frac{Q}{Q_{0}}\right)^{\nu}$. It follows that the de Sitter-Schwarzschild geometry acts as an attractor for this model, allowing us to define a boundary for the star and determine the mass function.
The structure of the paper is as follows: In Section II, we briefly introduce symmetric teleparallel gravity and its extension, the $f\left(Q\right)$-theory of gravity. Emphasis is given to the limit in which General Relativity is recovered within $f\left(Q\right)$-theory. In Section III, we focus on static spherically symmetric spacetimes. We introduce the symmetric and flat connection, which inherits the four symmetries of the background geometry, and present the gravitational field equations. We also discuss the de Sitter–Schwarzschild limit to be used for the exterior solution. In Section IV, we introduce the model under consideration; a power-law $f(Q)$-theory which offers a reasonable deviation from GR. Observational cosmological constraints on this theory have been studied in [66, 67]. We adopt a specific ansatz for the spacetime structure and analyze the evolution of the physical parameters of the gravitational fluid. We identify the radius of the star and, beyond this limit, we apply the necessary conditions so that the de Sitter–Schwarzschild solution describes the exterior. For the gravitational fluid of the interior solution, we derive the angular and radial speed of sound as well as the mass of the star. Finally, in Section V, we discuss our results and present our conclusions.
II Symmetric teleparallel gravity and modifications
In this section, we provide a brief overview of the fundamental aspects of the Symmetric Teleparallel Equivalent of General Relativity (STEGR) and its extensions. For a more detailed discussion, we refer the reader to [28, 68].
In metric-affine gravitational theories the space is characterized by two independent fields. The metric, $g_{\mu\nu}$, which serves as a rule of measuring distances, and the connection, $\Gamma_{~\mu\nu}^{\alpha}$, which is used to define the parallel transport of vectors. The most general connection is analyzed as follows:
$\Gamma_{~\mu\nu}^{\alpha}=\mathring{\Gamma}_{~\mu\nu}^{\alpha}+K_{~\mu\nu}^{\alpha}+L_{~\mu\nu}^{\alpha},$ (1)
where
$\mathring{\Gamma}_{~\mu\nu}^{\alpha}=\frac{1}{2}g^{\kappa\alpha}\left(\frac{\partial}{\partial x^{\mu}}g_{\kappa\nu}+\frac{\partial}{\partial x^{\nu}}g_{\mu\kappa}-\frac{\partial}{\partial x^{\kappa}}g_{\mu\nu}\right),$ (2)
denote the Christoffel symbols,
$K_{~\mu\nu}^{\alpha}=\frac{1}{2}\left(\mathcal{T}_{~\mu\nu}^{\alpha}+\mathcal{T}_{\mu~\nu}^{~a}+\mathcal{T}_{\nu~\mu}^{~a}\right),$ (3)
is the contorsion tensor defined with the help of the torsion, $\mathcal{T}_{~\mu\nu}^{\alpha}=2\left(\Gamma_{~\mu\nu}^{\alpha}-\Gamma_{~\nu\mu}^{\alpha}\right)$, and finally
$L_{~\mu\nu}^{\lambda}=\frac{1}{2}g^{\lambda\sigma}\left(Q_{\mu\nu\sigma}+Q_{\nu\mu\sigma}-Q_{\sigma\mu\nu}\right),$ (4)
is the disformation tensor constructed with the use of the nonmetricity
$Q_{\alpha\mu\nu}=\nabla_{\alpha}g_{\mu\nu}.$ (5)
In metric theories of gravity, like GR, the torsion and the nonmetricity tensors are zero. As a result, the connection coefficients are given in terms of the Christoffel symbols and thus the $\Gamma^{\alpha}_{~\mu\nu}$ is dependent on the metric. In metric-affine theories however, either one of the two - or even both - the torsion and the nonmetricity, are nonvanishing. In the case of $f(Q)$ gravity, only the nonmetricity is present. The connection is symmetric, ergo the torsion vanishes, and the spaces is taken to be flat, thus the $\Gamma_{~\mu\nu}^{\alpha}$ need to be such that the Riemann tensor
$R^{\lambda}_{~\mu\nu\kappa}=\frac{\partial}{\partial x^{\nu}}\Gamma_{~\mu\kappa}^{\lambda}-\frac{\partial}{\partial x^{\kappa}}\Gamma_{~\mu\nu}^{\lambda}+\Gamma_{~\mu\kappa}^{\eta}\Gamma_{~\nu\eta}^{\lambda}-\Gamma_{~\mu\nu}^{\eta}\Gamma_{~\kappa\eta}^{\lambda}$ (6)
is zero. In such a theory the gravitational effects are purely attributed to the nonmetricity.
The fundamental geometric object that enters in the action of the gravitational theories of this type is the nonmetricity scalar
$Q=-g^{\mu\nu}\left(L_{~\beta\mu}^{\alpha}L_{~\nu\alpha}^{\beta}-L_{~\beta\alpha}^{\alpha}L_{~\mu\nu}^{\beta}\right).$ (7)
The latter is constructed in such a way, so that the linear in $Q$ action
$S_{STEGR}=\int\!\!d^{4}x\sqrt{-g}Q,$ (8)
gives rise to field equations equivalent to those of GR; hence the name STEGR. This occurs because the nonmetricity $Q$ and the Ricci scalar of GR (the scalar derived from the curvature tensor constructed with the use of the Christoffel symbols), differ by a total divergence. In the following section we concentrate in the non-linear modifications of this theory.
II.1 $f\left(Q\right)$-theory
As happens in the case of GR and its modifications in terms of different $f(R)$ theories, we also obtain modifications to the STEGR by considering nonlinear functions of the fundamental scalar. We write the extensions of this theory by invoking the action
$S=\int\!\!d^{4}x\sqrt{-g}f(Q)+S_{m},$ (9)
where $f(Q)$ is generally a nonlinear function of $Q$. The $S_{m}$ denotes the part of the action corresponding to the matter content. For completeness, we need to mention that the full action includes the curvature $R^{\kappa}_{~\lambda\mu\nu}$ and the torsion tensor, which are added with Lagrange multipliers, so that the $R^{\kappa}_{~\lambda\mu\nu}=0$ and the $\mathcal{T}_{~\mu\nu}^{\alpha}=0$ conditions are induced as constraints [69, 70].
Variation of (9) with respect to the metric yields
$\frac{2}{\sqrt{-g}}\nabla_{\lambda}\left(\sqrt{-g}f_{Q}(Q)P_{\;\mu\nu}^{\lambda}\right)-\frac{1}{2}f(Q)g_{\mu\nu}+f_{Q}(Q)\left(P_{\mu\rho\sigma}Q_{\nu}^{\;\rho\sigma}-2Q_{\rho\sigma\mu}P_{~~~~\nu}^{\rho\sigma}\right)=\kappa\,T_{\mu\nu},$ (10)
where $\kappa=8\pi G/c^{4}$,
$P^{\alpha}_{\phantom{\alpha}\mu\nu}=-\frac{1}{4}Q^{\alpha}_{\phantom{\alpha}\mu\nu}+\frac{1}{2}Q_{(\mu\nu)}^{\phantom{(\mu\nu)}\alpha}+\frac{1}{4}\left(Q^{\alpha}-\tilde{Q}^{\alpha}\right)g_{\mu\nu}-\frac{1}{4}\delta^{\alpha}_{\phantom{\alpha}(\mu}Q_{\nu)}$ (11)
and $Q_{\alpha}=Q_{\alpha\phantom{\mu}\mu}^{\phantom{\alpha}\mu}$, $\tilde{Q}_{\alpha}=Q^{\mu}_{\phantom{\mu}\alpha\mu}$. The lower index $Q$ appearing in front of $f$ is used to symbolize the derivative of the function with respect to $Q$, i.e. $f_{Q}(Q)=\frac{df}{dQ}$, $f_{QQ}(Q)=\frac{d^{2}f}{dQ^{2}}$. The $T_{\mu\nu}$ is the energy-momentum tensor associated with the matter content which we add in the theory. Variation with respect to the (symmetric and flat) connection, which is an independent field from the metric, leads to the equations
$\nabla_{\mu}\nabla_{\nu}\left(\sqrt{-g}f_{Q}(Q)P^{\mu\nu}_{\phantom{\mu\nu}\alpha}\right)=0,$ (12)
assuming of course that the matter does not couple with the connection. Equation (12) is on equal footing with the field equations of the metric (10), it also needs to be satisfied. Not any connection for which (10) form a compatible set of equations will trivialize (12).
The field equations (10) of the metric can also be brought to the form [71]
$E_{\mu\nu}:=f_{Q}\left(Q\right)G_{\mu\nu}-\frac{1}{2}g_{\mu\nu}\left(f\left(Q\right)-f_{Q}\left(Q\right)Q\right)+2f_{QQ}\left(Q\right)P_{~~\mu\nu}^{\lambda}\nabla_{\lambda}Q=\kappa\,T_{\mu\nu},$ (13)
where $G_{\mu\nu}$ is the Einstein tensor of GR and $\kappa=8\pi G/c^{4}$. We use the $E_{\mu\nu}$ here to denote the components of the tensor expressing the gravitational part of the theory. A key observation is that, in contrast to what happens in GR, where $\nabla_{\mu}G^{\mu\nu}\equiv 0$, the divergence $\nabla_{\mu}E^{\mu\nu}$ is not identically zero. In order to have $\nabla_{\mu}E^{\mu\nu}=0$ the field equation for the connection, that is Eq. (12), needs to be satisfied [71]. This is a crucial point for the analysis that follows, because the conservation condition on the matter $\nabla_{\mu}T^{\mu\nu}=0$, creates a constraint among the functions of the metric, unless Eq. (12) has been satisfied before.
With the equations expressed in the form of (13) it is easy to see that the linear version of the theory simply reduces to $G_{\mu\nu}=\kappa_{\text{eff}}\;T_{\mu\nu}$. What is more, if $Q=$ const., then the above equations assume the form
$G_{\mu\nu}+\Lambda_{\text{eff}}\;g_{\mu\nu}=\kappa_{\text{eff}}\;T_{\mu\nu},$ (14)
where
$\Lambda_{\text{eff}}$ $=\frac{1}{2}\left(Q-\frac{f(Q)}{f_{Q}(Q)}\right)$ (15)
$\kappa_{\text{eff}}$ $=\frac{\kappa}{f_{Q}(Q)}$ (16)
are effective versions of the cosmological and Newton’s constants respectively. Thus, in the $f(Q)\sim Q$ or $Q=$ const. cases the theory is, at least at the level of the equations, equivalent to GR or GR plus cosmological constant respectively. Even though the gravitational $f(Q)$-theory is distinct from GR since gravity is now owed to nonmetricity, it is practically indistinguishable - at least at the level of dynamics - from GR. This is true even when considering the motion of matter. For example, a free particle just follows the extremal length curves (geodesics), which are identical irrespectively of whether the theory is based on nonmetricity or curvature. The geodesics just extremize $ds=\sqrt{-g_{\mu\nu}dx^{\mu}dx^{\nu}}$, which yields the usual equations involving the Christoffel symbols and not the generic connection.
III Static, Spherically Symmetric Spacetimes
III.1 Relativistic stars within General Relativity
When seeking interior solutions ($0\leq r\leq R$, $R$ being the radius of the star), the most general form of a static, spherically symmetric geometry in Schwarzschild-like coordinates, $\{t,r,\theta,\phi\}$, is given by
$ds^{2}=-e^{A(r)}dt^{2}+e^{B(r)}dr^{2}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right).$ (17)
with $A(r),B(r)$ being two independent functions of the radial coordinate. In the following we shall be working with the mass function, $m(r)$, which is defined by
$e^{-B}=1-\frac{2m}{r}.$ (18)
Anisotropic matter is described by a diagonal energy-momentum tensor of the form
$T^{\mu}_{\;\;\nu}=\mathrm{diag}\left(-\rho(r),p_{r}(r),p_{t}(r),p_{t}(r)\right),$ (19)
where $p_{r},p_{t}$ are the radial and tangential pressure of matter, respectively, $\rho$ is the energy density of the fluid, while the anisotropic factor is defined by
$\Delta=p_{t}-p_{r},$ (20)
while in the case of stars made of isotropic matter there is a unique pressure in both directions, $p_{r}=p_{t}$, and therefore $\Delta=0$.
In order to find solutions describing hydrostatic equilibrium of compact objects we have to integrate the Tolman-Oppenheimer-Volkoff equations [72, 73]
$m^{\prime}(r)$ $=$ $4\pi r^{2}\rho(r)$ (21)
$p_{r}^{\prime}(r)$ $=$ $-[\rho(r)+p_{r}(r)]\;\frac{A^{\prime}(r)}{2}+\frac{2\Delta}{r}$ (22)
$A^{\prime}(r)$ $=$ $2\>\frac{m(r)+4\pi r^{3}p_{r}(r)}{r^{2}\left(1-2m(r)/r\right)}$ (23)
where us usual a prime denotes differentiation with respect to $r$.
We integrate the structure equations throughout the star imposing at the center the following conditions
$m(0)=0,\;\;\;\;\;p_{r}(0)=p_{t}(0)=p_{c},$ (24)
where $p_{c}$ is the common central pressure. The anisotropic factor vanishes at the center of the star, $\Delta(0)=0$. We impose the following matching conditions at the surface of the star
$p_{r}(R)=0,\;\;\;\;m(R)=M,\;\;\;\;e^{A(R)}=1-2\frac{M}{R},$ (25)
with $M$ being the stellar mass, taking into account that the exterior vacuum solution ($r>R$) is given by the Schwarzschild geometry [57]
$ds^{2}=-(1-2M/r)dt^{2}+(1-2M/r)^{-1}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\>\phi^{2}).$ (26)
Finally, upon numerical integration, once the mass function and the radial pressure are known, the metric potential $A(r)$ is computed by
$A(r)=\ln\left(1-\frac{2M}{R}\right)+2\int^{r}_{R}dx\>\frac{m(x)+4\pi x^{3}p_{r}(x)}{x^{2}\left(1-2m(x)/x\right)}.$ (27)
III.2 Exact analytic solution within $f(Q)$ gravity
The theory is characterized by two basic fields, the metric and the connection. For the former we adopt the static and spherically symmetric line element as seen in Eq. (17). In what regards the connection, it is well known [74] that for a flat and symmetric connection there can always be found a coordinate system where its components are zero, i.e. $\Gamma_{~\mu\nu}^{\alpha}=0$. This is referred to the literature as the coincident gauge. However, we need to be careful when we impose such a condition. When assuming a particular class of metrics, like the one we read from (17), we need to consider that we have already fixed the gauge by choosing a particular coordinate system. The latter may happen to be incompatible with having $\Gamma_{~\mu\nu}^{\alpha}=0$.
In these cases, a guide in order to write connections which are compatible with the field equations, is given by requiring that the connection shares the symmetries of the metric, i.e. the isometries. In other words we demand that the Killing fields of the metric, satisfying $\mathcal{L}_{\xi}g_{\mu\nu}=0$, where $\mathcal{L}$ is the Lie-derivative, leave invariant the connection $\mathcal{L}_{\xi}\Gamma_{~\mu\nu}^{\alpha}=0$ as well. The resulting connections can play an important role in the dynamics introducing additional degrees of freedom [37]. For static, spherically symmetric spacetimes, there are two general families of connections with this property [75, 47]. The first leads to off-diagonal components in the field equations (10) for the gravitational part. These off-diagonal terms can only be eliminated by setting Q=Q=const. or $f(Q)\sim Q$. Alternatively, one can keep the off-diagonal terms and consider a non-diagonal energy momentum tensor for the matter. For the second family of connections the relative non-diagonal terms can be removed without requiring a constraint on the theory $f(Q)$ or on $Q$ itself. It is this second connection which we choose in this work. Its non-zero components are [75, 47, 50]:
$\begin{split}&\Gamma_{\;tt}^{t}=c_{1}+c_{2}-c_{1}c_{2}\gamma_{1},\quad\Gamma_{\;tr}^{t}=\frac{c_{2}\gamma_{1}(c_{1}\gamma_{1}-1)}{\gamma_{2}},\quad\Gamma_{\;rr}^{t}=\frac{\gamma_{1}(1-c_{1}\gamma_{1})(c_{2}\gamma_{1}+1)-\gamma_{2}\gamma_{1}^{\prime}}{\gamma_{2}^{2}},\\ &\Gamma_{\;\theta\theta}^{t}=-\gamma_{1},\quad\Gamma_{\phi\phi}^{t}=-\sin^{2}\theta\gamma_{1},\quad\Gamma_{\;tt}^{r}=-c_{1}c_{2}\gamma_{2},\quad\Gamma_{\;tr}^{r}=c_{1}\left(c_{2}\gamma_{1}+1\right),\\ &\Gamma_{\;rr}^{r}=\frac{1-c_{1}\gamma_{1}(c_{2}\gamma_{1}+2)-\gamma_{2}^{\prime}}{\gamma_{2}},\quad\Gamma_{\;\theta\theta}^{r}=-\gamma_{2},\quad\Gamma_{\;\phi\phi}^{r}=-\gamma_{2}\sin^{2}\theta,\quad\Gamma_{\;t\theta}^{\theta}=c_{1},\\ &\Gamma_{\;r\theta}^{\theta}=\frac{1-c_{1}\gamma_{1}}{\gamma_{2}},\quad\Gamma_{\phi\phi}^{\theta}=-\sin\theta\cos\theta,\quad\Gamma_{t\phi}^{\phi}=c_{1},\quad\Gamma_{\;r\phi}^{\phi}=\frac{1-c_{1}\gamma_{1}}{\gamma_{2}},\quad\Gamma_{\;\theta\phi}^{\phi}=\cot\theta,\end{split}$ (28)
where $c_{1}$, $c_{2}$ are constants and $\gamma_{1}(r)$, $\gamma_{2}(r)$ functions of the radial variable $r$. We use the prime to denote derivation with respect to $r$.
Non-metricity theories are based on two basic fields, the metric $g_{\mu\nu}$ and the connection $\Gamma_{\;\mu\nu}^{\kappa}$. The set of components (28) forms the initial ansatz for the connection $\Gamma_{\;\mu\nu}^{\kappa}$ in the same manner line element (17) constitutes the ansatz for the metric. For this choice of a connection, and for metric (17), the off-diagonal terms of the field equations (10) are eliminated if we further choose $c_{1}=c_{2}=0$.
By assuming a matter content leading to an anisotropic energy momentum tensor of the form (19), the remaining independent field equations for the metric read (for simplicity we drop the argument $Q$ from $f$ and its derivatives):
$\frac{e^{-B}}{r}\left[2\left(1-e^{\frac{B}{2}}\right)Q^{\prime}f_{QQ}-\frac{1}{r}\left(rB^{\prime}+e^{B}-1\right)f_{Q}\right]+\frac{1}{2}\left(Qf_{Q}-f\right)=\kappa\,T^{0}_{~0},$ (29a)
$\frac{1}{2}\left(Qf_{Q}-f\right)-\frac{e^{-B}}{r^{2}}\left(e^{B}-rA^{\prime}-1\right)f_{Q}=\kappa\,T^{1}_{~1},$ (29b)
$\frac{1}{2}\left(Qf_{Q}-f\right)+\frac{e^{-B}}{4r}\left[2rA^{\prime\prime}+\left(rA^{\prime}+2\right)\left(A^{\prime}-B^{\prime}\right)\right]f_{Q}-\frac{e^{-B}}{2r}\left(2e^{\frac{B}{2}}-rA^{\prime}-2\right)Q^{\prime}f_{QQ}=\kappa\,T^{2}_{~2}.$ (29c)
The equation (12) for the connection reduces to
$\begin{split}&Q^{\prime}f_{QQ}\left(A^{\prime}\left(r^{2}-e^{B}\gamma_{2}^{2}\right)-\left(e^{B}\gamma_{2}^{2}+r^{2}\right)B^{\prime}+4\left(r-e^{B}\gamma_{2}\gamma_{2}^{\prime}\right)\right)+2\left(r^{2}-e^{B}\gamma_{2}^{2}\right)Q^{\prime 2}f_{QQQ}\\ &+2\left(r^{2}-e^{B}\gamma_{2}^{2}\right)Q^{\prime\prime}f_{QQ}=0,\end{split}$ (30)
while the nonmetricity scalar (7) is expressed as
$Q=\frac{e^{-B}}{r^{2}\gamma_{2}^{2}}\left[2r^{2}\gamma_{2}^{\prime}-e^{B}\gamma_{2}^{3}\left(A^{\prime}+B^{\prime}\right)+r\gamma_{2}\left(-rA^{\prime}+rB^{\prime}-4\right)+2\gamma_{2}^{2}\left(rA^{\prime}-e^{B}\gamma_{2}^{\prime}+e^{B}+1\right)\right].$ (31)
A key observation here is that the simple choice
$\gamma_{2}(r)=re^{-\frac{B}{2}}$ (32)
satisfies equation (30) irrespectively of the $f(Q)$ function and of the $A(r)$ and $B(r)$ appearing in the metric. This connection has been also utilized in the context of scalar non-metricity to derive black hole solutions [22].
A brief comment is in order at this point. Many works in the literature, attempting to study stellar configurations in $f(Q)$ gravity, either incorporate an overrestricting procedure or include the serious error of disregarding the equation of the connection. In some cases, the coincident gauge is adopted, which however is incompatible with line element (17) in spherical coordinates and thus overrestricts the equations; leading only to solutions of the linear $f(Q)\sim Q$ theory, which is however equivalent to GR [76, 77, 78, 79, 80]. In other cases, even though a connection compatible with spherical symmetry is considered, it is usually not checked whether the equation of motion for the connection (12) is satisfied. To make this problem more explicit; a common choice of a connection used in the literature [81, 82, 83, 84, 85] is the following
$\begin{split}&\Gamma^{r}_{\;\theta\theta}=-r,\quad\Gamma^{r}_{\;\phi\phi}=-r\sin^{2}\theta,\quad\Gamma^{\theta}_{\;r\theta}=\Gamma^{\theta}_{\;\theta r}=\Gamma^{\phi}_{\;r\phi}=\Gamma^{\phi}_{\;\phi r}=\frac{1}{r},\quad\Gamma^{\theta}_{\;\phi\phi}=-\sin\theta\cos\theta,\\ &\Gamma^{\phi}_{\;\theta\phi}=\Gamma^{\phi}_{\;\phi\theta}=\cot\theta,\end{split}$ (33)
with the rest of components being zero. The resulting nonmetricity scalar is now
$Q=\frac{e^{-B}\left(1-e^{B}\right)\left(A^{\prime}+B^{\prime}\right)}{r}.$ (34)
It can be seen that connection (33) is a special case of (28) with $c_{1}=c_{2}=\gamma_{1}=0$ and $\gamma_{2}(r)=r$. The latter choice for $\gamma_{2}$ however, unlike (32), does not satisfy the field equation for the connection (30), which in this case yields a restriction between the functions $A$ and $B$ of the metric
$\left[\sinh\left(\frac{B}{2}\right)\left(\left(rA^{\prime}+4\right)Q^{\prime}+2rQ^{\prime\prime}\right)+rB^{\prime}\cosh\left(\frac{B}{2}\right)Q^{\prime}\right]f_{QQ}+2r\sinh\left(\frac{B}{2}\right)Q^{\prime 2}f_{QQQ}=0.$ (35)
Remember that $Q$ here is given by (34), and for a given $f(Q)$ theory this is going to be purely a relation between the functions of the metric, $A$ and $B$. The above equation is trivialized if Q=Q=const. or $f(Q)\sim Q$, which, as is known, describe the same solutions as in the case of GR with or without cosmological constant. Thus, in the generic setting of a non-trivial $f(Q)$ theory, it is not sufficient for the description of the star solution to just solve algebraically (29) with respect to the energy density and the pressures. The constraint set by the field equation of the connection needs to be satisfied. Note that it is not enough to satisfy the analogue of the Tolman-Oppenheimer-Volkoff (TOV) equation. The $\nabla_{\mu}T^{\mu\nu}=0$ condition on the matter content together with the field equations for the metric (29) are not sufficient conditions. This is because, as we previously mentioned, the covariant derivative of the gravitational tensor $E_{\mu\nu}$ of Eq. (13), unlike what happens in GR, is not identically zero. The $\nabla_{\mu}E^{\mu\nu}=0$ equation holds only if the equation of the connection (30) is satisfied [71]. In order for the latter to happen for a theory dynamically different from GR, we need to find an appropriate $\gamma_{2}(r)$ function that solves the connection equation.
In our case, the $\gamma_{2}(r)$ we derived in (32) serves to satisfy the field equation for the connection (30), and thus the constraint in the metric functions is overall avoided. What is more, (32) satisfies the relevant equation for any $f(Q)$ theory. In this respect, the $\gamma_{2,\text{int}}=re^{-\frac{B}{2}}$ forms an excellent choice for the connection in what regards the interior solution for the star. For the exterior vacuum solution we adopt a different path, which we describe in detail in the next section. It consists of adopting an exterior solution, $\gamma_{2,\text{ext}}$, for which the non-metricity scalar becomes constant. Thus, we achieve for the inside, a consistent description of modified gravitational dynamics due to a dynamical non-metricity, while in the exterior a known vacuum solution compatible with GR plus a cosmological constant. In the same manner that one connects two distinct solutions for the interior (with matter) and for the exterior (vacuum) for the metric, we additionally connect here the two distinct solutions for the other basic field of the theory, the connection.
III.3 Exterior solution
We know that the theory becomes dynamically equivalent to GR plus a cosmological constant when Q=Q=const. Hence, the equations are going to be satisfied by the Schwarzschild-de Sitter metric if, outside the star and in vacuum, we take the nonmetricity scalar to be constant. As a result, for the exterior region, we choose to use
$ds^{2}_{\text{ext}}=-\left(1-\frac{2M}{r}-\frac{\Lambda_{\text{eff}}}{3}r^{2}\right)dt^{2}+\frac{dr^{2}}{\left(1-\frac{2M}{r}-\frac{\Lambda_{\text{eff}}}{3}r^{2}\right)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta d\phi^{2}\right),$ (36)
where the effective cosmological constant $\Lambda_{\text{eff}}$ is given by (15).
In order for $Q$ to be constant, we need $\gamma_{2}$ to satisfy the differential equation $Q=q$, $q\in\mathbb{R}$, where $Q$ is given by (31). This equation can be easily integrated to give
$\begin{split}\gamma_{2}=&\gamma_{2,\text{ext}}=\frac{1}{12f_{Q}}\Bigg[r^{3}f-C_{1}-2r\left(q\,r^{2}-6\right)f_{Q}\\ &\pm\left[\left(C_{1}-r^{3}f\right)^{2}-4r\left(q\,r^{5}f-C_{1}qr^{2}+6C_{1}\right)f_{Q}+4r\left(72M+q\,r^{3}\left(q\,r^{2}-6\right)\right)f_{Q}^{2}\right]^{\frac{1}{2}}\Bigg],\end{split}$ (37)
with $C_{1}$ being an integration constant and $f_{Q}=f_{Q}(q)=$const. The constant $C_{1}$ plays a crucial role since with this we can continuously connect the $\gamma_{2,\text{ext}}$ with the $\gamma_{2,\text{int}}$ given by (32) at the radius of the star. The equation of motion for the connection now is satisfied due to the fact that the nonmetricity $Q$ is a constant.
IV The model under consideration
Let us consider a theory of the form
$f(Q)=Q+\alpha\,Q_{0}\left(\frac{Q}{Q_{0}}\right)^{\nu},$ (38)
where $\alpha$, $Q_{0}$ and $\nu$ are constants. The ansatz we follow for the metric functions $A$ and $B$ is the following [86]
$A(r)$ $=C+\left(\frac{r}{r_{1}}\right)^{2},$ (39)
$B(r)$ $=\left(\frac{r}{r_{2}}\right)^{2},$ (40)
with $C\in\mathbb{R}$ and $r_{1}$, $r_{2}$ positive constants. The corresponding geometry describes a singular-free physical solution.
With these considerations, the resulting energy density and the two pressures from (29) become
$\begin{split}\rho=&\frac{Q_{0}^{-\nu}e^{-\frac{r^{2}}{r_{2}^{2}}}}{2\kappa r^{2}Q^{2}}\Bigg[4\alpha(\nu-1)\nu Q_{0}r\left(e^{\frac{r^{2}}{2r_{2}^{2}}}-1\right)Q^{\nu}Q^{\prime}+2Q^{2}Q_{0}^{\nu}\left(\frac{2r^{2}}{r_{2}^{2}}+e^{\frac{r^{2}}{r_{2}^{2}}}-1\right)\\ &+2\alpha\nu Q_{0}\left(\frac{2r^{2}}{r_{2}^{2}}+e^{\frac{r^{2}}{r_{2}^{2}}}-1\right)Q^{\nu+1}-\alpha(\nu-1)Q_{0}r^{2}e^{\frac{r^{2}}{r_{2}^{2}}}Q^{\nu+2}\Bigg],\end{split}$ (41)
$\begin{split}p_{r}=\frac{Q_{0}^{-\nu}e^{-\frac{r^{2}}{r_{2}^{2}}}}{2\kappa r^{2}r_{1}^{2}Q}&\Bigg[\alpha Q_{0}Q^{\nu}\left(r_{1}^{2}e^{\frac{r^{2}}{r_{2}^{2}}}\left((\nu-1)r^{2}Q-2\nu\right)+2\nu\left(2r^{2}+r_{1}^{2}\right)\right)\\ &+2QQ_{0}^{\nu}\left(2r^{2}-r_{1}^{2}\left(e^{\frac{r^{2}}{r_{2}^{2}}}-1\right)\right)\Bigg],\end{split}$ (42)
$\begin{split}p_{t}=\frac{Q_{0}^{-\nu}e^{-\frac{r^{2}}{r_{2}^{2}}}}{2\kappa rr_{1}^{4}r_{2}^{2}Q^{2}}&\Bigg[\alpha Q_{0}Q^{\nu}\left((2(\nu-1)\nu)r_{1}^{2}r_{2}^{2}Q^{\prime}\left(r^{2}-r_{1}^{2}\left(e^{\frac{r^{2}}{2r_{2}^{2}}}-1\right)\right)+(\nu-1)rr_{1}^{4}r_{2}^{2}Q^{2}e^{\frac{r^{2}}{r_{2}^{2}}}\right)\\ &-2rQ\left(r^{2}\left(r_{1}^{2}-r_{2}^{2}\right)+r_{1}^{4}-2r_{1}^{2}r_{2}^{2}\right)\left(QQ_{0}^{\nu}+\alpha\nu Q_{0}Q^{\nu}\right)\Bigg],\end{split}$ (43)
respectively. The nonmetricity scalar appearing in the above relations is calculated from (31), with the use of (32) for the interior, which leads to
$Q(r)=\frac{2e^{-B}}{r^{2}}\left(e^{\frac{B}{2}}-1\right)\left(-rA^{\prime}+e^{\frac{B}{2}}-1\right)=\frac{2e^{-\frac{r^{2}}{r_{2}^{2}}}}{r^{2}}\left(e^{\frac{r^{2}}{2r_{2}^{2}}}-1\right)\left(e^{\frac{r^{2}}{2r_{2}^{2}}}-\frac{2r^{2}}{r_{1}^{2}}-1\right).$ (44)
Using Eq. (44) in the expressions (41)-(43), and considering theories $f(Q)$ for which $\nu>1$ in (38), it is straightforward to demonstrate that the relevant physical quantities assume a finite value at $r=0$. The nonmetricity scalar itself becomes zero at $r\rightarrow 0$. In the Taylor expansion, there always exists a zero-th order term, while the rest that follow are in positive powers of $r$. Specifically, for $\nu\geq 2$ the leading $r$-dependent terms are quadratic:
$\rho$ $=\frac{3}{\kappa r_{2}^{2}}+\mathcal{O}(r^{2})$ (45)
$p_{r}$ $=\frac{1}{\kappa}\left(\frac{2}{r_{1}^{2}}-\frac{1}{r_{2}^{2}}\right)+\mathcal{O}(r^{2})$ (46)
$p_{t}$ $=\frac{1}{\kappa}\left(\frac{2}{r_{1}^{2}}-\frac{1}{r_{2}^{2}}\right)+\mathcal{O}(r^{2}).$ (47)
In (45)-(47) we distinguish the finite values at the origin for the energy density and the two pressures. The pressures are equal at $r=0$, but differentiate as $r>0$.
Turning to calculate the speed of sound in the radial and tangential directions from
$c_{s,\text{radial}}^{2}=\frac{p_{r}^{\prime}}{\rho^{\prime}},\quad c_{s,\text{tangential}}^{2}=\frac{p_{t}^{\prime}}{\rho^{\prime}}$ (48)
we obtain at the limits $r\rightarrow 0$
$\lim\limits_{r\to 0}c_{s,\text{radial}}^{2}=\frac{\left(4r_{2}^{2}-r_{1}^{2}\right)\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+4\alpha r_{2}^{2}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)},$ (49)
$\lim\limits_{r\to 0}c_{s,\text{tangential}}^{2}=\frac{4\alpha\left(r_{1}^{4}-6r_{1}^{2}r_{2}^{2}+8r_{2}^{4}\right)-2Q_{0}r_{2}^{2}\left(r_{1}^{4}-3r_{1}^{2}r_{2}^{2}+r_{2}^{4}\right)}{5r_{1}^{2}\left(Q_{0}r_{1}^{2}r_{2}^{2}-2\alpha r_{1}^{2}+8\alpha r_{2}^{2}\right)}$ (50)
for a theory with $\nu=2$, while for $\nu$ assuming natural number values with $\nu>2$ we recover
$\lim\limits_{r\to 0}c_{s,\text{radial}}^{2}=\frac{4r_{2}^{2}}{5r_{1}^{2}}-\frac{1}{5},$ (51)
$\lim\limits_{r\to 0}c_{s,\text{tangential}}^{2}=\frac{2\left(3r_{1}^{2}r_{2}^{2}-r_{1}^{4}-r_{2}^{4}\right)}{5r_{1}^{4}}.$ (52)
In Fig. 1 we present the plots of the two speeds of sound in the interior of the star for certain values of the parameters. In particular, we have assumed $r_{1}=r_{2}=10\>km$, $Q_{0}=1/r_{1}^{2}$ and $\alpha=0.01,0.1$. The values of the parameters $r_{1},r_{2}$ are representative, as they are comparable to the stellar radius.
The radius of the star is defined as the distance $r=\mathcal{R}$, where $p_{r}(\mathcal{R})=0$. As can be seen from (42), this condition defines a transcendental equation, which can only be solved numerically for certain values of the parameters. We calculate the mass, $m(r)$, of the star by parameterizing the $g_{rr}$ component of the metric as
$\left(1-\frac{2m(r)}{r}-\frac{\Lambda_{\text{eff}}}{3}r^{2}\right)^{-1}=e^{\frac{r^{2}}{r_{2}^{2}}},$ (53)
with
$\Lambda_{\text{eff}}=\frac{1}{2}\left(Q(\mathcal{R})-\frac{f(Q(\mathcal{R}))}{f_{Q}(Q(\mathcal{R}))}\right).$ (54)
At the radius, $r=\mathcal{R}$, the nonmetricity assumes a value $q=Q(\mathcal{R})$, which we consider to remain constant in the outside region $r>r_{s}$. The resulting mass function for $r\leq\mathcal{R}$ from (53) is
$m(r)=\frac{1}{2}r\left(1-\frac{\Lambda_{\text{eff}}}{3}r^{2}-e^{-\frac{r^{2}}{r_{2}^{2}}}\right).$ (55)
The total mass of the star is of course given by $m(\mathcal{R})=M$. In Fig. 2 we depict the function of the mass divided by the mass of the Sun ($M_{\text{Sun}}\sim 1.989\times 10^{30}$ kg) as a function of the normalized radial distance $r/\mathcal{R}$. As can be seen, stronger deviations from GR, lead to more rapidly rising curves with increasing radius.
Figure 1: The radial (continuous lines) and tangential (dashed lines) sound speeds as functions of the star radius. The theory considered has $\nu=2$. The red lines correspond to $\alpha=0.1$ and the blue ones to $\alpha=0.01$. For the rest of the parameters we have $r_{1}=r_{2}=10$ km and $Q_{0}=0.01$ km-2.
Figure 2: The mass function in the interior of the star. The power index characterizing the theory is $\nu=2$. The red lines correspond to $\alpha=0.1$ and the blue ones to $\alpha=0.01$. For both curves $r_{1}=r_{2}=10$ km and $Q_{0}=0.01$ km-2.
Figure 3: Normalized energy density (dotted curves) and pressures in both directions (radial: solid curves, tangential: dashed curves) versus radial coordinate assuming $\alpha=0.01$ in blue color and $\alpha=0.1$ in red color. All three quantities are positive throughout the star, while at the same time $\rho>p_{r,t}$.
Figure 4: Relativistic adiabatic index, $\Gamma$, in both directions (radial: solid, tangential: dashed) versus radial coordinate considering $\alpha=0.01$ in blue color and $\alpha=0.1$ in red color. The horizontal line in black represents the Newtonian value $4/3$.
For the sake of comparison, we report here the stellar mass and radius for the cases of GR ($\alpha=0$) as well as the two numerical values for $\alpha=0.01,0.1$ for the $f(Q)$ gravity considered here. They are as follows:
$R=11.21km,\;\;\;\;\;M=2.70M_{\odot},\;\;\;\;\;\alpha=0$ (56)
$R=11.24km,\;\;\;\;\;M=2.71M_{\odot},\;\;\;\;\;\alpha=0.01$ (57)
$R=11.53km,\;\;\;\;\;M=2.81M_{\odot},\;\;\;\;\;\alpha=0.1$ (58)
Finally, in order to further demonstrate that the solution found here is capable of describing realistic astronomical configurations, we show in Fig. 3 and Fig. 4 that the energy conditions [87, 88] are fulfilled, and that the stability criterion based on the relativistic adiabatic index, $\Gamma>4/3$, is met, with $4/3$ corresponding to the Newtonian value [89, 90]. In particular, for compact stars made of anisotropic matter the energy conditions are the following, see e.g. [91]:
$\rho\geq 0,\;\;\;\;\;\rho+p_{r}+2p_{t}\geq 0,\;\;\;\;\;\rho\pm p_{i}\geq 0,\;\;\;i=r,t$ (59)
while the relativistic adiabatic index, in both directions, is defined by
$\Gamma_{r}=c_{s,r}^{2}\left(1+\frac{\rho}{p_{r}}\right),\;\;\;\;\;\Gamma_{t}=c_{s,t}^{2}\left(1+\frac{\rho}{p_{t}}\right).$ (60)
It is observed that both $\Gamma_{r},\Gamma_{t}$ remain higher than $4/3$, while at the same time the energy density remains larger than both pressures. Moreover, both pressures and the energy density are positive throughout the star. Therefore, the stability criterion based on $\Gamma$ is met, and all energy conditions are fulfilled.
V Conclusions
In this study, we investigated the structure of static spherically symmetric stars within the framework of symmetric teleparallel $f\left(Q\right)$-gravity. The limit of General Relativity (GR), with or without the cosmological constant, is recovered in $f\left(Q\right)$-theory when $Q$ is constant. Consequently, if $Q$ is a non-zero constant, the de Sitter–Schwarzschild solution emerges.
We exploited the existence of this known vacuum solution to use it as the exterior metric in our problem. For the interior of the star, we encountered first the connection which is compatible with the field equations induced by a space-time given by line element (17). It is important to note that this connection can be used in this context with any $f(Q)$ theory. In our investigation we used a power-law $f(Q)$ function (38) and, under the ansatz (39), we determined the physical properties of the respective fluid. We determined variables of the parameters, which give a reasonable behavior for the energy density and the two pressures. We identified the radius $r=\mathcal{R}$, where we assume that $Q(\mathcal{R})=\text{const}$. for $r\geq\mathcal{R}$. In this setting we calculated the explicit expression for the mass function.
Additionally, we analyzed the radial and tangential sound speeds near the singularity. For values $\nu>2$ of the power index characterizing the $f\left(Q\right)$-theory, the variation of the sound speeds is nearly independent of $\nu$. These findings contribute to a deeper understanding of the role of nonmetricity in gravitational models and support the viability of $f(Q)$ gravity in describing compact objects. In a future work we plan to study the definition for the Tolman - Oppenheimer - Volkov formula within the symmetric teleparallel theory of gravity.
Acknowledgements.
AG was supported by Proyecto Fondecyt Regular 1240247. AP thanks the support of Vicerrectoría de Investigación y Desarrollo Tecnológico (Vridt) at Universidad Católica del Norte through Núcleo de Investigación Geometría Diferencial y Aplicaciones, Resolución Vridt No - 096/2022 and Resolución Vridt No - 098/2022. AP was partially supported by Proyecto Fondecyt Regular 2024, Folio 1240514, Etapa 2024. AP thanks the Universidad de La Frontera for the hospitality provided when this work was carried out.